# Bridging doubly heavy tetraquark mass spectrum with heavy baryons utilizing heavy antiquark-diquark symmetry

### v1

@@ III Results and discussions
Table 5: Spectrum of DHTs aroused from heavy anti-baryons $\bar{\Omega}_{c}$($\bar{\Omega}_{b}$) with HADS in this work and comparisons with different models or methods. DMC: diffusion Monte Carlo method. CMI: color-magnetic interaction model; RQM(VM): relativized quark model within the variational method; RQM(DA): relativized quark model within diquark-antidiquark picture; MIT: MIT bag model; ECM: extended chromomagnetic model.
Configuration $J^{P}$ This work DMC Mutuk (2024) CMI Luo et al. (2017) MIT Zhang et al. (2021) RQM(VM) Lü et al. (2020) RQM(DA) Ebert et al. (2007) ECM I Weng et al. (2022) ECM II Weng et al. (2022)
$\bar{s}\bar{s}]^{3}_{1}[cc]^{\bar{3}}_{1}]_{0}$ $0^{+}$ $T_{cc\bar{s}\bar{s}}(4224)$ $4205$ $4395$ $4300$ $4417$ $4359$ $4043.7$ $4199.1$
$\bar{s}\bar{s}]^{3}_{1}[cc]^{\bar{3}}_{1}]_{1}$ $1^{+}$ $T_{cc\bar{s}\bar{s}}(4239)$ $4323$ $4526$ $4382$ $4493$ $4375$ $4192.6$ $4300.2$
$\bar{s}\bar{s}]^{3}_{1}[cc]^{\bar{3}}_{1}]_{2}$ $2^{+}$ $T_{cc\bar{s}\bar{s}}(4270)$ $4381$ $4597$ $4433$ $4536$ $4402$ $4264.5$ $4372.1$
$\bar{s}\bar{s}]^{3}_{1}[bb]^{\bar{3}}_{1}]_{0}$ $0^{+}$ $T_{bb\bar{s}\bar{s}}(10740)$ $10868$ $11157$ $11078$ $10972$ $10932$ $10697.1$ $10792.1$
$\bar{s}\bar{s}]^{3}_{1}[bb]^{\bar{3}}_{1}]_{1}$ $1^{+}$ $T_{bb\bar{s}\bar{s}}(10745)$ $10908$ $11199$ $11099$ $10986$ $10939$ $10718.2$ $10809.8$
$\bar{s}\bar{s}]^{3}_{1}[bb]^{\bar{3}}_{1}]_{2}$ $2^{+}$ $T_{bb\bar{s}\bar{s}}(10755)$ $10926$ $11224$ $11119$ $11004$ $10950$ $10742.5$ $10834.1$
$\bar{s}\bar{s}]^{3}_{1}[cb]^{\bar{3}}_{0}]_{1}$ $1^{+}$ $T_{cb\bar{s}\bar{s}}(7508)$ $7711$ $7793$ $7716$ $7752$ $7683$ $7431.8$ $7553.6$
$\bar{s}\bar{s}]^{3}_{1}[cb]^{\bar{3}}_{1}]_{0}$ $0^{+}$ $T_{cb\bar{s}\bar{s}}(7527)$ $7653$ $7774$ $7693$ $7735$ $7673$ $7404.4$ $7531.3$
$\bar{s}\bar{s}]^{3}_{1}[cb]^{\bar{3}}_{1}]_{1}$ $1^{+}$ $T_{cb\bar{s}\bar{s}}(7535)$ $7764$ $7872$ $7757$ $7775$ $7684$ $7503.3$ $7603.5$
$\bar{s}\bar{s}]^{3}_{1}[cb]^{\bar{3}}_{1}]_{2}$ $2^{+}$ $T_{cb\bar{s}\bar{s}}(7550)$ $7862$ $7908$ $7779$ $7798$ $7701$ $7534.3$ $7633.8$
@@
Common characteristics include the universal mass hierarchy $T_{QQ^{\prime}}<T_{QQ^{\prime}\bar{s}}<T_{QQ^{\prime}\bar{s}\bar{s}}$, reduced hyperfine splittings in heavier flavor sectors, and enhanced binding for $b$-quark systems due to diminished kinetic energy. The progressive stabilization from $cc$ to $bb$ systems highlights the mass-dependence of binding mechanisms, while systematic threshold comparisons provide crucial diagnostics for distinguishing compact tetraquarks from molecular configurations.
@@ IV Summary
In this work, we provides a comprehensive investigation of DHTs ($T_{QQ\bar{q}\bar{q}^{\prime}}$) spectrum using HADS within a well accepted constituent quark model framework calibrated to established hadron data. By correlating $QQ\bar{q}\bar{q}^{\prime}$ states with experimentally measured heavy baryon ($Qqq^{\prime}$) spectrum through HADS, we systematically predict masses for 38 ground-state tetraquarks with $J^{P}=0^{+},1^{+},2^{+}$ across $cc$, bbbb, and $bc$ flavor sectors with varying strange quark content. Key findings reveal that $bb$-containing tetraquarks predominantly lie below meson-meson thresholds, indicating strong binding and potential stability, while $cc$ and $bc$ configurations generally reside near or above thresholds except for specific cases like the $T_{cc}(3876)$ state that matches the observed $T_{cc}^{+}(3875)$. The universal mass ordering $T_{QQ^{\prime}}<T_{QQ^{\prime}\bar{s}}<T_{QQ^{\prime}\bar{s}\bar{s}}$ and method-independent hyperfine splittings provide critical discriminants between compact tetraquarks and molecular interpretations. All parameters derive from baryon spectra fits, ensuring model consistency with QCD phenomenology without new inputs. These predictions establish essential benchmarks for experimental searches at LHCb, Belle II, and PANDA, while demonstrating HADS as a fundamental symmetry unifying conventional and exotic heavy-quark hadrons.

### v2

@@ III Results and discussions
Table 10: Spectrum of DHTs aroused from heavy anti-baryons $\bar{\Omega}_{c}$($\bar{\Omega}_{b}$) with HADS (values in brackets are without HADS) in this work and comparisons with different models or methods. DMC: diffusion Monte Carlo method. CMI: color-magnetic interaction model; RQM(VM): relativized quark model within the variational method; RQM(DA): relativized quark model within diquark-antidiquark picture; MIT: MIT bag model; ECM: extended chromomagnetic model.
$J^{P}$ Configuration This work DMC Mutuk (2024) CMI Luo et al. (2017) MIT Zhang et al. (2021) RQM(VM)  et al. (2020) RQM(DA) Ebert et al. (2007) ECM I Weng et al. (2022) ECM II Weng et al. (2022)
$0^{+}$ $\bar{s}\bar{s}]^{3}_{1}[cc]^{\bar{3}}_{1}]_{0}$ $4224$ $4205$ $4395$ $4300$ $4417$ $4359$ $4043.7$ $4199.1$
$1^{+}$ $\bar{s}\bar{s}]^{3}_{1}[cc]^{\bar{3}}_{1}]_{1}$ $4239$ $4323$ $4526$ $4382$ $4493$ $4375$ $4192.6$ $4300.2$
$2^{+}$ $\bar{s}\bar{s}]^{3}_{1}[cc]^{\bar{3}}_{1}]_{2}$ $4270$ $4381$ $4597$ $4433$ $4536$ $4402$ $4264.5$ $4372.1$
$0^{+}$ $\bar{s}\bar{s}]^{3}_{1}[bb]^{\bar{3}}_{1}]_{0}$ $10740$ $10868$ $11157$ $11078$ $10972$ $10932$ $10697.1$ $10792.1$
$1^{+}$ $\bar{s}\bar{s}]^{3}_{1}[bb]^{\bar{3}}_{1}]_{1}$ $10745$ $10908$ $11199$ $11099$ $10986$ $10939$ $10718.2$ $10809.8$
$2^{+}$ $\bar{s}\bar{s}]^{3}_{1}[bb]^{\bar{3}}_{1}]_{2}$ $10755$ $10926$ $11224$ $11119$ $11004$ $10950$ $10742.5$ $10834.1$
$0^{+}$ $\bar{s}\bar{s}]^{3}_{1}[cb]^{\bar{3}}_{1}]_{0}$ $7527$ $7653$ $7774$ $7693$ $7735$ $7673$ $7404.4$ $7531.3$
$1^{+}$ $\bar{s}\bar{s}]^{3}_{1}[cb]^{\bar{3}}_{0}]_{1}$ $7508(7499)$ $7711$ $7793$ $7716$ $7752$ $7683$ $7431.8$ $7553.6$
$\bar{s}\bar{s}]^{3}_{1}[cb]^{\bar{3}}_{1}]_{1}$ $7535(7531)$ $7764$ $7872$ $7757$ $7775$ $7684$ $7503.3$ $7603.5$
$2^{+}$ $\bar{s}\bar{s}]^{3}_{1}[cb]^{\bar{3}}_{1}]_{2}$ $7550$ $7862$ $7908$ $7779$ $7798$ $7701$ $7534.3$ $7633.8$
@@
Common characteristics include the universal mass hierarchy $T_{QQ^{\prime}}<T_{QQ^{\prime}\bar{s}}<T_{QQ^{\prime}\bar{s}\bar{s}}$, reduced hyperfine splittings in heavier flavor sectors, and enhanced binding for $b$-quark systems due to diminished kinetic energy. The progressive stabilization from $cc$ to $bb$ systems highlights the mass-dependence of binding mechanisms. Furthermore, by systematically comparing the obtained mass spectrum of DHTs with the corresponding meson-meson thresholds, we find that most of the DHT masses are above or far from the thresholds. Considering that molecular states are generally close to and below the thresholds, this provides crucial diagnostics for distinguishing compact tetraquarks from corresponding molecular states.
@@ IV Summary
In this work, we provides a comprehensive investigation of DHTs ($T_{QQ\bar{q}\bar{q}^{\prime}}$) spectrum using HADS within a well accepted constituent quark model framework calibrated to established hadron data. By correlating $QQ\bar{q}\bar{q}^{\prime}$ states with experimentally measured heavy baryon ($Qqq^{\prime}$) spectrum through HADS, we systematically predict masses for 38 ground-state tetraquarks with $J^{P}=0^{+},1^{+},2^{+}$ across $cc$, bbbb, and $bc$ flavor sectors with varying strange quark content. All parameters derive from baryon spectra fits, ensuring model consistency with QCD phenomenology without new inputs. Key findings reveal that $bb$-containing tetraquarks predominantly lie below meson-meson thresholds, indicating strong binding and potential stability, while $cc$ and $bc$ configurations generally reside near or above thresholds except for specific cases like the $T_{cc}(3876)$ state that matches the observed $T_{cc}^{+}(3875)$. The universal mass ordering $T_{QQ^{\prime}}<T_{QQ^{\prime}\bar{s}}<T_{QQ^{\prime}\bar{s}\bar{s}}$ and method-independent hyperfine splittings provide critical discriminants between compact tetraquarks and molecular interpretations.
Our study shows that the hyperfine splitting of DHT states decreases with the increase in the mass of heavy diquarks, which is consistent with the requirements of heavy quark symmetry. We have also estimated the impact of HADS symmetry breaking on the mass of DHTs, and found that for double-charmed, bottom-charmed, and double-bottomed tetraquark states, the impact levels are about 11.3 MeV, 5.7 MeV, and 3.8 MeV respectively. This order relationship also reflects the heavy quark mass dependence of heavy quark symmetry. These predictions establish essential benchmarks for experimental searches at LHCb, Belle II, and PANDA, while demonstrating HADS as a fundamental symmetry unifying conventional and exotic heavy-quark hadrons.